John makes 20 drinks in 5 minutes. Steve makes 20 drinks in 10 minutes and Sue makes 20 drinks in 15 minutes. How long will it take to make 20 drinks if they worked together?
Answer by checkley77(12844)

1/5+1/10+1/15=1/X FIND THE LCD (30).
(6+3+2)/30=1/X
11/30=1/X CROSS MULTIPLY.
11X=30
X=30/11
X=2.727 MINUTES TO MIX 20 DRINKS WHEN ALL 3 WORK TOGETHER.